What is the smallest positive phase constant which is equivalent to 7.5pi

Grade:11

1 Answers

Saurabh Kumar
9 years ago
Phase constant is the sine of angle of the angular shift.
Calculate the sine of angular shift.
sin(2\pi + \pi + \pi/2) = c